OpenStudy (anonymous):

\[{d \over dx} \ln {e^x \over 1+e^x}\]

OpenStudy (anonymous):

d/dx(e^x/(1+e^x)/ [e^x/(1+e^x)] = [e^x(1+e^x)-e^x.(e^x) ] / [e^x/(1+e^x)]= 1+e^x

OpenStudy (anonymous):

the quotient rule and the derivative of ln

OpenStudy (lalaly):

\[\ln \frac{e^x}{1+e^x}=\ln(e^x)-\ln(1+e^x)\]

OpenStudy (lalaly):

\[=x-\ln(1+e^x)\]

OpenStudy (lalaly):

now its easy to differentiate

OpenStudy (anonymous):

so the answer would be 1- (e^x)/(1+e^x)?

OpenStudy (mimi_x3):

Yes, the answer is: \[1-\frac{e^{x}}{1+e^{x}} \]

OpenStudy (anonymous):

hmm, that's what I had at the beginning but i didn't get it on my calculator

OpenStudy (mimi_x3):

\[x-\ln(1+e^x) =>1-\frac{1}{1+e^{x}} *e^x = 1-\frac{e^{x}}{1+e^{x}} = \frac{1+e^{x}-e^{x}}{1+e^{x}} = \frac{1}{1+e^{x}} \]

OpenStudy (anonymous):

...oh. could've done that I guess :P Thanks, all!